AZHAR H. MIRZA

6207 Springhill Court Apt. 203 Greenbelt MD 20770 Tel/Fax: (301) 513-5939 (h) Tel: (301) 405-6572 (w) E-mail : azzaro@eng.umd.edu

PROFILE

STRENGTHS

Systems Engineer with strong background in Computer Integrated Manufacturing, Production Planning, Facility Design, System and Software Life Cycle Development, Client-Server development, Object Oriented Design, Analysis and Programming.

COMPUTER SKILLS

• Languages : C, C++, SQL, Pascal, Fortran.

• Operating Systems : UNIX (System calls, Network programming, IPC & Shell), MS-DOS, Mac.

• Platforms : SUN Workstation (SunOS, Solaris X-window), DEC Station 5000, IBM-PC, NeXT.

• Database Systems : INGRES, ObjectStore.

• GUI Systems : Tcl-Tk, SUIT (Xlib/Motif based GUI toolkit)

EDUCATION

M.S. in Mechanical Engineering UNIVERSITY OF MARYLAND, College Park Expected : 8/95

• Thesis : Computer Aided Design & Implementation of Hybrid Manufacturing Facility. GPA 3.4

B.S. in Mechanical Engineering KARNATAKA REGIONAL ENGINEERING COLLEGE, Surathkal 8/92

• Ranked 3rd out of 200 students of Mechanical Engineering at Mangalore University.

EXPERIENCE

Researcher & Software Developer 9/93 – Present

COMPUTER INTEGRATED MANUFACTURING LABORATORY, Institute for Systems Research, University of Maryland, College Park and WESTINGHOUSE ESG, Baltimore.

• Hybrid Facility Re-design Software : Designed, implemented and documented a manufacturing shop redesign software system in C/Unix with a SUIT GUI. Employed the developed software to design a hybrid shop layout and devised an optimal plan for its implementation resulting in a benefit of $1,460,000 on $452,000 investment for Westinghouse ESG.

Planning Engineer 8/92 – 8/93

MOTOR INDUSTRIES COMPANY LTD, India (subsidiary Robert Bosch GmbH)

• Rotational placements in Production Management, Process Planning and MIS departments. Designed and developed PC-based software tools for inventory monitoring & control and computer aided process planning, using C, dBase IV and Lotus-123.

TECHNICAL PROJECTS

• Production Simulation : Simulated the production operation of a fortune 500 company using software coded in C++ employing principles of discrete event dynamic systems. The resulting solution was compared with the results of the analytical model and simulation of the manufacturing facility using SIMFACTORY simulation software.

• Hierarchical Production Planning : Designed and implemented a hierarchical production planning software system in C to determine the optimal workforce, production, overtime and sub-contracting levels. The resulting plan was compared with the plan obtained from the MRP II system (ROI MANAGE 2000).

• Client Server Architecture, Network Programming & Unix System V IPC : Developed the client

server architecture for an Object Oriented, hypertext capable, Document Annotation Facility, using BSD socket and XDR. Developed concurrency control and deadlock prevention system using Message Queues and a database of system resources using Shared Memory.

• Relational and Object Oriented Database Design : Designed and developed a database application

for management of media rental store using embedded SQL in C on INGRES and C++ on ObjectStore. The relative performance characteristics of relational and object oriented databases were evaluated and compared.

R end help h q quit end